OpenStudy (anonymous):

Write the equation in standard form. Identify the important features of the graph: x^2 + y^2 - 8x + 10y + 15 = 0

OpenStudy (anonymous):

\[ x^2 + y^2 - 8x + 10y + 15 = 0\\ x^2 - 8x +16 -16 + y^2 +10y + 25 - 25+ 15 =0\\ (x-4)^2 -16 + (y+5)^2 -25+ 15 =0\\ (x-4)^2 + (y+5)^2 =-15+25+16=26\ \] Can you finish it now?

OpenStudy (kc_kennylau):

@eliassaab I thought you've done it?